# HEPHY-ML-25-02 BitHEP — The Limits of Low-Precision ML in HEP Claudius Krause1, Daohan Wang1 and Ramon Winterhalder2

### v1

@@ 2.1 BitLinear layer
As illustrated in Fig. 1, both layers begin by mapping the trainable weights onto a quantized representation . The choice of quantization depends on the acceptable trade-off between precision and efficiency. The standard approach [162] uses binary weights, , while an extended version [163] allows for ternary weights, , improving feature filtering. Both weight quantizations are parametrized as
| binary (1-bit): | |||||
| ternary (1.58-bit): | (1) |
@@
2.2 Implementation and code availabilty
@@ 3 Classification: Quark-gluon tagging
Jet clustering is performed using the anti- algorithm with a radius parameter of using FastJet [173]. We select only jets with transverse momentum [500, 550] GeV and rapidity for further analysis. The dataset contains not only the four-momenta of each particle but also particle identification labels, including electron, muon, charged hadron, neutral hadron, and photon.
@@ 3.1 Particle Dual Attention Transformer
Specifically, we employ QAT for the particle and channel attention modules while keeping other components, including the feature extractor, 1D CNN, and final multi-layer perceptron (MLP) classifier, in full precision. This is achieved by replacing all linear layers with BitLinear layers in the two particle attention and two channel attention modules, affecting approximately 63% of the total weight parameters. The attention modules incorporate physics-motivated bias terms in the scaled dot-product attention, which remain in full precision to preserve critical information. The final output undergoes a 1D CNN transformation and global average pooling before being fed into an MLP classifier. Further architectural details can be found in Ref. [76]. The new model is trained from scratch, with all hyperparameters identical to the non-quantized version.
@@ 3.2 Performance comparison
As depicted in Tab. 1, while P-DAT itself offers robust performance with an accuracy of 0.839 and an AUC of 0.9092, the adaptation of this model into P-DAT-Bit reveals both the strengths and the trade-offs of utilizing BitLinear layers within the attention blocks. The integration of BitLinear layers in P-DAT-Bit results in a slight decrease in accuracy (0.834) and AUC (0.9040) compared to its predecessor. Despite this modest drop, the performance metrics remain highly competitive, indicating that the reduced precision computation does not drastically compromise the model’s discriminative capability. Specifically, the background rejection rates at 50% signal efficiency (Rej50%) and 30% signal efficiency (Rej30%) are 35.0 and 83.3, respectively. This variation between P-DAT and P-DAT-Bit highlights the balancing act between computational efficiency and performance accuracy. It emphasizes that while BitLinear layers streamline model operations—particularly advantageous in resource-constrained environments—there is a nuanced impact on the model’s ability to manage complex discriminative tasks.
@@
Finally, in Fig. 2, we present the calibration curves of P-DAT and P-DAT-Bit models for quark/gluon discrimination. The red line represents the calibration curve for the original P-DAT model, while the blue line represents the P-DAT-Bit model. Each model was evaluated over five runs to compute the average calibration curve and the associated standard deviation. Both models demonstrate good calibration performance, with their curves closely following the diagonal line, indicating that the predicted probabilities are well-calibrated with the actual positive class probabilities. Notably, the error bands for both models are very small, suggesting that the predictions of both models are consistent across multiple runs. Besides, the blue band of the P-DAT-Bit model is slightly broader compared to the red band of the P-DAT model, which can be attributed to its use of lower bit precision for weights and inputs in 60% of its parameters. Despite this, the P-DAT-Bit model maintains a good calibration performance, closely following the ideal diagonal line, indicating that its predicted probabilities are generally accurate.
@@ 4.1 SMEFTNet architecture
SMEFTNet is an IRC-safe and rotation-equivariant graph neural network. It is designed to provide an optimal observable for small deviations from the SM and enhance SMEFT sensitivity, specifically focusing on the linear SM-SMEFT interference within SMEFT. To preserve sensitivity to the linear term, it is crucial to consider the orientation of the decay planes of the or boson, as this orientation helps resolve the helicity configuration of the amplitude that is altered in the SMEFT [178]. To study the hadronic final states of or boson, SMEFTNet is constructed to be equivariant to azimuthal rotations of the constituents of the boosted jet around the jet axis, maintaining SO(2) symmetry regardless of the chosen reference frame. It processes inputs as variable-length lists of particle constituents of a fat jet, originating from the hadronic decay of a boosted massive particle. For details on the architecture, we refer to Ref. [117].
In our study, we consider three quantized variants of the SMEFTNet model and apply the three variants to predict the decay plane angle from the jet’s constituents and compare their performance with the results presented in Ref [117]. In the first variant, all linear layers are replaced with BitLinear layers, which we refer to as SMEFTNet-Bit. In the second variant, only the linear layers within the MLP block are quantized, corresponding to approximately 70% of the total weight parameters; this configuration is labeled SMEFTNet-Bit70. In the third variant, only the linear layers in the Message Passing Neural Network block are quantized, accounting for about 30% of the weights; this is denoted SMEFTNet-Bit30.
@@ 4.2 Performance comparison
To visualize the comparative analysis of SMEFTNet and SMEFTNet-Bit, Fig. 3 presents the scatter plots of the true and the predicted in the remaining 20% dataset, respectively. The results reveal that while the SMEFTNet-Bit model marginally underperforms relative to the original SMEFTNet, the differences are minimal, showcasing the effectiveness of the low-bit model in capturing the essential structure of the data despite its limited weight and input precision. Notably, the scatter plot of SMEFTNet exhibits two faint vertical red bands at true values of , which become more pronounced and broader in the SMEFTNet-Bit scatter plot. This phenomenon is attributed to both models’ tendency to stabilize at local minima due to the inherent challenges posed by multi-objective optimization. With two regression targets at the critical true values, each model confronts conflicting optimal solutions that necessitate a compromise, potentially deviating from the global optimum for any single objective. The broader bands in the SMEFTNet-Bit results are explained by its limited precision due to weight and input quantization. This restricted precision leads to significant fluctuations in the predicted values, resulting in the observed wider bands in the visual representations.
@@
Finally, Fig. 4 presents three histograms of residuals, which represent the differences between truths and predictions for . In each histogram, the red bars correspond to the original SMEFTNet, while the blue bars represent different configurations of the SMEFTNet-Bit model. In the left histogram, the blue bars correspond to SMEFTNet-Bit where all linear layers in SMEFTNet are replaced with BitNet layers. In the middle histogram, the blue bars represent the SMEFTNet-Bit70 where the linear layer weights in the MLP block are quantized, accounting for 70% of the total weight parameters. In the right histogram, the blue bars represent the SMEFTNet-Bit30 with quantized linear layer weights in the Message Passing NN block, accounting for 30% of the total weight parameters. In all cases, the SMEFTNet model exhibits a sharp peak at zero, indicating a precise alignment of predictions with true values. When fully quantized (100%), the SMEFTNet-Bit distribution broadens prominently, reflecting larger variations in prediction accuracy. Reducing the fraction of quantized layers to 70% yields intermediate performance: the residuals of SMEFTNet-Bit70 remain somewhat more dispersed than SMEFTNet but are considerably narrower than the fully quantized version. At 30% quantization, the residuals of SMEFTNet-Bit30 closely resemble those of the original SMEFTNet, showing only a modest increase in width. Consequently, it is evident that as more linear layers are replaced by BitNet layers, the performance of the model deteriorates. Moreover, a notable feature in all plots is the periodic structure of the residuals, which arises because the model faces a periodic ambiguity in which angles separated by map to the same predicted value, creating these peaks around and 0 in the residual distribution. Furthermore, the fraction at is higher for SMEFTNet-Bit, reflecting its broader vertical band in its 2D scatter plots. This broadening stems from the limited precision of BitLinear layers, which amplifies prediction fluctuations near critical values of . Notably, as the proportion of quantized layers increases, the fraction at also becomes higher, indicating that stronger quantization further destabilizes the predictions near critical values of . By contrast, SMEFTNet shows narrower bands, indicating more accurate predictions and a lower fraction at . Lastly, across all three configurations, both models overlap around zero residuals, highlighting their overall reliability. However, the progressive widening of SMEFTNet-Bit’s distributions with increasing quantization underscores the trade-off between model compression and predictive precision—demonstrating that partial, rather than complete, quantization can give a better balance for resource-limited applications.
@@ 5 Generative: Detector simulation
Next, we consider a generative task. These are very important in HEP, as the process of sampling random events from a given (complicated) conditional probability density is precisely what happens in numerical simulations of particle collisions (See [1] for on overview on ML applications to the entire simulation chain.). One notable challenge is detector simulations, which involve modeling showers of energetic particles in the calorimeters—a task characterized by its high numerical complexity and dimensionality. In the past years, this field has seen many new ideas and approaches [179], which led to the conception of the CaloChallenge [180] to compare existing models on equal footing and spur even more development of state-of-the-art methods. The CaloChallenge was a data challenge in the HEP community, with four different datasets, increasing in their dimensionality. The goal of the challenge was to train generative networks on the datasets and to generate artificial samples as fast and precise as possible [181]. The dataset dimensionalities range from a few hundred in the easiest case to a few tens of thousands in the most complicated case. These are single particle showers, simulated with Geant4 [182, 183, 184] and available at [185, 186, 187, 188]. For more details on the datasets we refer to [180].
As an example for generative networks, we consider two well-performing submissions of the CaloChallenge [180]—CaloINN [55] based on a normalizing flow and CaloDREAM [56] based on conditional flow matching. Normalizing flows were the first generative model that passed the “classifier test” [39, 189] in calorimeter shower simulations and have seen various applications to this task [57, 55, 190, 191, 192, 48, 39, 41, 46, 180]. Conditional flow matching is the most recent generative model that was explored in this context and shows impressive performance in many applications [56, 193, 194], possibly becoming more widely used in the future.
5.1 Normalizing Flow — CaloINN
CaloINN [55] is a normalizing-flow-based generative model, learning a bijective transformation between a simple base distribution—usually a Gaussian—and a more complicated distribution—the target distribution. In detail, CaloINN is based on coupling layer-based normalizing flow [195, 196] (which are sometimes also called INN, invertible neural networks [197], even though they all are invertible). These types of flows are equally fast to evaluate in both directions (as density estimator and generative model), making training and generation more efficient than in an autoregressive setup [39, 41, 46]. The bijective transformation that is used in the coupling blocks is based on splines [198, 199, 200].

### v2

@@ 2.1 BitLinear layer
As illustrated in Fig. 1, both layers begin by mapping the trainable weights onto a quantized representation q. The choice of quantization depends on the acceptable trade-off between precision and efficiency. The standard approach [162] uses binary weights, , while an extended version [163] allows for ternary weights, , improving feature filtering. Both weight quantizations are parametrized as
| binary (1-bit): | q | ||||
| ternary (1.58-bit): | q | (1) |
@@
2.2 Implementation and code availability
@@
2.3 Computational resource requirements
The relative timing of different numerical operations, e.g. floating-point operations (FLOPs) or integer operations (IntOPs), is highly hardware dependent. Since we want to factorize out this component from our analysis, we report on the number of required operations instead. This will serve as a proxy for the resource complexity. The number of operations of a regular linear layer, including a bias term but excluding an activation function, is
where is the batch size and denotes the size of the input/output. In the BitLinear and BitLinear158 layer, the number of operations is
The “less than” factor arises because zero weights in the ternary quantization do not contribute to the sum, so the number of IntOPs is smaller than the sum over all elements. The three FLOPs come from the scaling with , , and the bias term. The term independent of the output size comes from the initial quantization of the input vector, which is performed once for all outputs.
The timing comparisons among FLOP, SignOP, and IntOP are hardware dependent, with dedicated hardware obviously much faster than the general-purpose hardware usually available. In general, the SignOP is basically free, as it is a copy/bitflip/skip operation for weights on any hardware. The IntOP depends on the kind of FLOP one compares against and the kind of integer used. Here, we have integer additions, which are even cheaper than integer multiplications. An estimated difference between this IntOP and a FLOP is of the order of a factor 10-30.
Figure 2 shows the ratio of the two scalings for the conservative assumption 1 FLOP = 10 IntOPs and 1 FLOP = 30 IntOPs, for different input and output sizes. In all cases, we assume vanishing costs for SignOPs and the maximal number of IntOPs (i.e. no zero weights that reduce the size of the sum).
Based on these considerations, it is clear that two networks with the same number of trainable parameters and the same fraction of quantization can still achieve very different speed-ups depending on the number of layers and their widths within the network.
@@ 3 Classification: Quark-gluon tagging
Jet clustering is performed using the anti- algorithm with a radius parameter of using FastJet [173]. We select only jets with transverse momentum [500, 550] GeV and rapidity for further analysis. The dataset contains not only the four-momenta of each particle but also particle identification labels, including electron, muon, photon, as well as individual categories for the different charged and neutral hadrons, i.e. the full PID parameterization is used.
@@ 3.1 Particle Dual Attention Transformer
Specifically, we employ QAT for the particle and channel attention modules while keeping other components, including the feature extractor, 1D CNN, and final multi-layer perceptron (MLP) classifier, in full precision. This choice is motivated by computational efficiency and robustness considerations. The attention modules contain the majority (6̃3) of parameters and are the most computationally demanding, thus offering the greatest benefit from quantization. Furthermore, transformer-based attention mechanisms have been shown to exhibit notable robustness to quantization, making them an ideal starting point for a proof-of-concept study. Therefore, we implement QAT by replacing all linear layers with BitLinear layers in the two particle attention and two channel attention modules, affecting approximately 63% of the total weight parameters. The attention modules incorporate physics-motivated bias terms in the scaled dot-product attention, which remain in full precision to preserve critical information. The final output undergoes a 1D CNN transformation and global average pooling before being fed into an MLP classifier. Further architectural details can be found in Ref. [76]. The new model is trained from scratch, with all hyperparameters identical to the non-quantized version.
@@ 3.2 Performance comparison
As depicted in Tab. 1, while P-DAT itself offers robust performance with an accuracy of 0.839 and an AUC of 0.9092, the adaptation of this model into P-DAT-Bit reveals both the strengths and the trade-offs of utilizing BitLinear layers within the attention blocks. The integration of BitLinear layers in P-DAT-Bit results in a slight decrease in accuracy and AUC compared to the non-quantized variant. Despite this modest drop, the performance metrics remain highly competitive, indicating that the reduced precision computation does not drastically compromise the model’s discriminative capability. Specifically, the background rejection rates at 50% signal efficiency (Rej50%) and 30% signal efficiency (Rej30%) are 35.0 and 83.3, respectively. This variation between P-DAT and P-DAT-Bit highlights the balancing act between computational efficiency and performance accuracy. It emphasizes that while BitLinear layers streamline model operations,particularly advantageous in resource-constrained environments, there is a nuanced impact on the model’s ability to manage complex discriminative tasks.
Finally, in Fig. 3, we present the calibration curves of P-DAT and P-DAT-Bit models for quark/gluon discrimination. The red line represents the calibration curve for the original P-DAT model, while the blue line represents the P-DAT-Bit model. Each model was evaluated over five runs to compute the average calibration curve and the associated standard deviation. Both models demonstrate good calibration performance, with their curves closely following the diagonal line, indicating that the predicted probabilities are well-calibrated with the actual positive class probabilities. Notably, the error bands for both models are very small, suggesting that the predictions of both models are consistent across multiple runs. Besides, the blue band of the P-DAT-Bit model is slightly broader compared to the red band of the P-DAT model, which can be attributed to its use of lower bit precision for weights and inputs in 60% of its parameters. Despite this, the P-DAT-Bit model is well calibrated and closely following the ideal diagonal line, indicating that its predicted probabilities are generally accurate.
To quantify the computational cost of the P-DAT-Bit, we compute the number of FLOPs, IntOPs, and SignOPs of its two building blocks, the Particle Attention Block and the Channel Attention Block in Tab. 2.
| Block | Operations | P-DAT | P-DAT-Bit | |
| Particle Attention Block | FLOPs | |||
| SignOPs | - | |||
| IntOPs | - | |||
| Total FLOPs | (41%) | |||
| Channel Attention Block | FLOPs | |||
| SignOPs | - | |||
| IntOPs | - | |||
| Total FLOPs | (34%) | |||
@@ 4.1 SMEFTNet architecture
SMEFTNet is an IRC-safe and rotation-equivariant graph neural network. It is designed to provide an optimal observable for small deviations from the SM and enhance SMEFT sensitivity, specifically focusing on the linear SM-SMEFT interference within SMEFT. To preserve sensitivity to the linear term, it is crucial to consider the orientation of the decay planes of the or boson, as this orientation helps resolve the helicity configuration of the amplitude that is altered in the SMEFT [178]. To study the hadronic final states of or boson, SMEFTNet is constructed to be equivariant to azimuthal rotations of the constituents of the boosted jet around the jet axis, maintaining SO(2) symmetry regardless of the chosen reference frame. It processes inputs as variable-length lists of particle constituents of a fat jet, originating from the hadronic decay of a boosted massive particle. For details on the architecture, we refer to Ref. [117]. Although the decay-plane angle regression is a relatively lightweight task, investigating quantization in this controlled setting is still meaningful. In realistic SMEFT inference tasks, surrogate models have to be evaluated repeatedly when profiling over many nuisance parameters to estimate systematic uncertainties, or when integrating over latent observables, where lighter and faster models become highly advantageous, and even moderate speedups can lead to substantial reductions in end-to-end inference time. Thus, assessing whether low-precision variants of lightweight models retain accuracy is a necessary first step before applying such techniques in more computationally intensive inference workflows.
In our study, we consider three quantized variants of the SMEFTNet model and apply the three variants to predict the decay plane angle from the jet’s constituents and compare their performance with the results presented in Ref [117]. In the first variant, all linear layers are replaced with BitLinear layers, which we refer to as SMEFTNet-Bit100. In the second variant, only the linear layers within the MLP block are quantized, corresponding to approximately 70% of the total weight parameters; this configuration is labeled SMEFTNet-Bit70. In the third variant, only the linear layers in the Message Passing Neural Network block are quantized, accounting for about 30% of the weights; this is denoted SMEFTNet-Bit30.
@@ 4.2 Performance comparison
| Model Comparison | Wasserstein Distance | Separation Power |
| SMEFTNet | 0.0021 | 0.0001 |
| SMEFTNet-Bit100 | 0.4546 | 1.4860 |
| SMEFTNet-Bit70 | 0.2528 | 0.6138 |
| SMEFTNet-Bit30 | 0.1040 | 0.1503 |
@@
To visualize the comparative analysis of SMEFTNet and SMEFTNet-Bit100, Fig. 4 presents the scatter plots of the true and the predicted in the remaining 20% dataset, respectively. The results reveal that while the SMEFTNet-Bit100 model marginally underperforms relative to the original SMEFTNet, the differences are minimal, showcasing the effectiveness of the low-bit model in capturing the essential structure of the data despite its limited weight and input precision. Notably, the scatter plot of SMEFTNet exhibits two faint vertical red bands at true values of , which become more pronounced and broader in the SMEFTNet-Bit100 scatter plot. This phenomenon is attributed to both models’ tendency to stabilize at local minima due to the inherent challenges posed by multi-objective optimization. In the critical region of true , two physically indistinguishable parton-level configurations — arising from the permutation of up-type and down-type quarks — correspond to decay plane angles that differ by . Since the particle-level information does not resolve this ambiguity, the regression model is trained with both values as plausible targets. As a result, it is forced to compromise between conflicting objectives, potentially leading to suboptimal predictions for either configuration individually. The broader bands in the SMEFTNet-Bit100 results are explained by its limited precision due to weight and input quantization. This restricted precision leads to significant fluctuations in the predicted values, resulting in the observed wider bands in the visual representations.
@@
Finally, Fig. 5 presents three histograms of residuals, which represent the differences between truths and predictions for . In each histogram, the red bars correspond to the original SMEFTNet, while the blue bars represent different configurations of the SMEFTNet-Bit model. In the left histogram, the blue bars correspond to SMEFTNet-Bit100. In the middle histogram, the blue bars represent the SMEFTNet-Bit70. In the right histogram, the blue bars represent the SMEFTNet-Bit30. In all cases, the SMEFTNet model exhibits a sharp peak at zero, indicating a precise alignment of predictions with true values. When fully quantized (100%), the SMEFTNet-Bit100 distribution broadens prominently, reflecting larger variations in prediction accuracy. Reducing the fraction of quantized layers to 70% yields intermediate performance: the residuals of SMEFTNet-Bit70 remain somewhat more dispersed than SMEFTNet but are considerably narrower than the fully quantized version. At 30% quantization, the residuals of SMEFTNet-Bit30 closely resemble those of the original SMEFTNet, showing only a modest increase in width. Consequently, it is evident that as more linear layers are replaced by BitNet layers, the performance of the model deteriorates. Moreover, a notable feature in all plots is the periodic structure of the residuals, which arises because the model faces a periodic ambiguity in which angles separated by map to the same predicted value, creating these peaks around and 0 in the residual distribution. Furthermore, the fraction at is higher for SMEFTNet-Bit100, reflecting its broader vertical band in its 2D scatter plots. This broadening stems from the limited precision of BitLinear layers, which amplifies prediction fluctuations near critical values of . Notably, as the proportion of quantized layers increases, the fraction at also becomes higher, indicating that stronger quantization further destabilizes the predictions near critical values of . By contrast, SMEFTNet shows narrower bands, indicating more accurate predictions and a lower fraction at . Lastly, across all three configurations, both models overlap around zero residuals, highlighting their overall reliability. However, the progressive widening of SMEFTNet-Bit’s distributions with increasing quantization underscores the trade-off between model compression and predictive precision. This demonstrates that partial, rather than complete, quantization can give a better balance for resource-limited applications.
Table 3 complements Fig. 5 with two quantitative metrics. The Wasserstein distance captures the average angular shift between residual distributions, whereas the separation power reflects their overall shape divergence. The additional “SMEFTNet” entry corresponds to a second model trained with the same architecture and hyper-parameters as the reference model but initialized with a different random seed. Its near-zero Wasserstein distance (0.002) and separation power () serve as a self-consistency check, confirming that both metrics are sensitive only to distributional shifts and robust to statistical fluctuations. By contrast, the three quantized variants show progressively larger values as the proportion of BitLinear layers increases, quantitatively substantiating the broadening already visible in Fig. 5.
| Block | Operations | SMEFTNet | SMEFTNet-Bit100 | |
| Message Passing Block | FLOPs | |||
| SignOPs | - | |||
| IntOPs | - | |||
| Total FLOPs | (27%) | |||
| MLP Block | FLOPs | |||
| SignOPs | - | |||
| IntOPs | - | |||
| Total FLOPs | (20%) | |||
To estimate the resource consumption of the SMEFTNet-Bit100, we compute the number of FLOPs, IntOPs, and SignOPs of its two building blocks, the Message Passing Block and the MLP Block in Tab. 4. Since the Message Passing Block is applied once per edge while the MLP Block is applied once per event, we report their per-pass costs separately.
@@ 5 Generative: Detector simulation
Next, we consider a generative task. These are very important in HEP, as the process of sampling random events from a given (complicated) conditional probability density is precisely what happens in numerical simulations of particle collisions (See [1] for an overview on ML applications to the entire simulation chain.). One notable challenge is detector simulations, which involve modeling showers of energetic particles in the calorimeters, a task characterized by its high numerical complexity and dimensionality. In the past years, this field has seen many new ideas and approaches [179], which led to the conception of the CaloChallenge [180] to compare existing models on equal footing and spur even more development of state-of-the-art methods. The CaloChallenge was a data challenge in the HEP community, with four different datasets, increasing in their dimensionality. The goal of the challenge was to train generative networks on the datasets and to generate artificial samples as fast and precise as possible [181]. The dataset dimensionalities range from a few hundred in the easiest case to a few tens of thousands in the most complicated case. These are single particle showers, simulated with Geant4 [182, 183, 184] and available at [185, 186, 187, 188]. For more details on the datasets we refer to [180].
When evaluating generative models, one can pick from various metrics [189, 190, 180]. Here, we use classifiers to evaluate the generative performance. The classifiers are trained on the task of distinguishing generated samples from the Geant4 reference. If a powerful, well-trained classifier cannot distinguish between the two samples, we conclude that the samples were drawn from the same underlying distribution, i.e. the generative model learned the underlying distribution well [39]. Summarized in a single quantity, we use the area under the receiver operating characteristic (AUC). A lower AUC indicates a better generative model.
As an example for generative networks, we consider two well-performing submissions of the CaloChallenge [180]: CaloINN [55] based on a normalizing flow and CaloDREAM [56] based on conditional flow matching. Normalizing flows were the first generative model that passed the “classifier test” [39, 190] in calorimeter shower simulations and have seen various applications to this task [57, 55, 191, 192, 193, 48, 39, 41, 46, 180]. Conditional flow matching is the most recent generative model that was explored in this context and shows impressive performance in many applications [56, 194, 195].
5.1 Normalizing Flow – CaloINN
CaloINN [55] is a normalizing-flow-based generative model, learning a bijective transformation between a simple base distribution, usually a Gaussian, and a more complicated distribution, the target distribution. In detail, CaloINN is based on coupling layer-based normalizing flow [196, 197] (which are sometimes also called INN, invertible neural networks [198], even though they all are invertible). These types of flows are equally fast to evaluate in both directions (as density estimator and generative model), making training and generation more efficient than in an autoregressive setup [39, 41, 46]. The bijective transformation that is used in the coupling blocks is based on splines [199, 200, 201].